OpenStudy (anonymous):

HELP BUT TELL ME HOW TO DO NOT JUST AN ANSWER rewrite equation to make y a function of x x+2=1/3(y-4)

hero (hero):

Is it: \[x+2 = \frac{1}{3(y-4)}\] or \[x + 2 = \frac{(y-4)}{3}\] Believe it or not, with the way you wrote it, it can be interpreted more than one way. So just say which one is correct.

hero (hero):

Maybe it is printed in your document as \[x + 2 = \frac{1}{3}(y-4)\]

OpenStudy (anonymous):

yes that is how it is written in my document

OpenStudy (anonymous):

what i attached is supposed to be a walk through, that i don't get

hero (hero):

What is it exactly that you don't get?

hero (hero):

When they say to re-write y as a function of x, what they mean is to isolate y.

hero (hero):

In other words, to re-write the equation so that y is on one side by itself, and the expression for it is on the other side. In order to do so, you must manipulate the equation using algebraic operations.

hero (hero):

@nocontact112